OpenStudy (anonymous):

Solve this system of equations. x-z=5 y+3z=12 2x+y=7

OpenStudy (anonymous):

2x-2z-2x-y=5-7 -2z-y=-2 -2z-y+y+3z=12-2 z=10 x=5+z=15 y=12-3z=12-30=-18

OpenStudy (anonymous):

Thanks!

OpenStudy (campbell_st):

I think you need to check the calculations and solutions....as they don't work for the last equation 2x + y = 7 of x = 15 and y = -18 30-18 = 12.... and not 7